# The Angular Momentum Penrose Inequality

### v3

@@ 1.4 Relation to Prior Work
The Angular Momentum Penrose Inequality has a substantial history dating to Penrose’s original proposal [48]. Our work synthesizes existing methods—specifically the Jang equation approach of Bray–Khuri [11] and Han–Khuri [29], the area-angular momentum inequalities of Dain–Reiris [22], and the $p$-harmonic flow of Agostiniani–Mazzieri–Oronzio [1]—into a complete proof for the axisymmetric vacuum case.
Technical contributions:
  • AM-Hawking mass functional: We introduce the functional $m_{H,J}(t)=\sqrt{m_{H}^{2}+4\pi J^{2}/A(t)}$ and establish its monotonicity by combining Hawking mass monotonicity with the Dain–Reiris sub-extremality bound.
  • Angular momentum conservation: We prove that $J(\Sigma_{t})$ is constant along the $p$-harmonic level set flow under the vacuum hypothesis (Theorem 6.14).
  • Axisymmetric Jang equation: We extend the Han–Khuri analysis to incorporate twist potentials, establishing the necessary decay estimates for the twist perturbation near the MOTS (Theorem 4.12).
Relation to prior work:
  • Time-symmetric Penrose inequality (Huisken–Ilmanen [32], Bray [10]): These foundational works established the $J=0$ case using IMCF and conformal flow respectively. Our result extends to $J\neq 0$ using different flow methods (AMO $p$-harmonic flow).
  • Spacetime Penrose inequality (Bray–Khuri [11], Han–Khuri [29]): The Jang equation approach and the scalar curvature identity $R_{\bar{g}}\geq 2|\boldsymbol{j}|^{2}$ are due to Bray–Khuri. Han–Khuri developed the MOTS blow-up analysis that we rely upon.
  • Area-angular momentum inequalities (Dain [20], Dain–Reiris [22]): The fundamental bound $A\geq 8\pi|J|$ for stable MOTS is their result. We use this as an input—we do not reprove it. Our contribution is showing how this bound integrates into the mass inequality proof.
  • AMO flows [1]: The $p$-harmonic flow and Hawking mass monotonicity are established by Agostiniani–Mazzieri–Oronzio. We adapt their framework to track angular momentum.
  • Mars–Simon uniqueness [83, 84]: The Simon tensor characterization of Kerr is classical. We apply these methods in our specific geometric setting.
Remark 1.16 (Initial Data Result).
Theorem 1.2 is a statement about initial data—a Riemannian 3-manifold $(M,g)$ with symmetric 2-tensor $K$ satisfying the constraint equations. It does not require or use any information about the future time evolution of this data. The inequality is proven using geometric analysis on the fixed initial data slice, not dynamical arguments.
@@ 1.5 Organization
The paper is organized as follows. In Section 2, we verify the inequality for the Kerr spacetime. Section 3 outlines the proof strategy. Section 4 establishes the existence of solutions to the axisymmetric Jang equation. Section 5 treats the angular-momentum-modified Lichnerowicz equation. Section 6 details the AMO flow and angular momentum conservation. Section 7 incorporates the sub-extremality bound. Section 8 synthesizes these results to prove the main theorem. Finally, Section 9 addresses the rigidity case, and Section 10 discusses extensions.
@@ 2 Verification for Kerr Spacetime
We first verify that the Kerr solution saturates the inequality with equality.
Remark 2.1 (Purpose of This Section).
Verifying that the conjectured equality case (Kerr) actually saturates the bound is a necessary consistency check for any Penrose-type inequality. If Kerr failed to saturate the bound, the conjecture would be wrong. This verification also determines the correct functional form of the bound—specifically, the combination $A/(16\pi)+4\pi J^{2}/A$ appearing in (2).
Definition 2.2 (Kerr Parameters).
We work in geometric units ($G=c=1$). For the Kerr spacetime with mass $M$ and spin parameter $a=J/M$ (where $|a|\leq M$ for sub-extremality):
$M_{\mathrm{ADM}}$ $=M,$ (4)
$J$ $=aM,$ (5)
$r_{+}$ $=M+\sqrt{M^{2}-a^{2}}\quad\text{(outer horizon radius)},$ (6)
$A$ $=4\pi(r_{+}^{2}+a^{2})=8\pi M(M+\sqrt{M^{2}-a^{2}})\quad\text{(horizon area)}.$ (7)
@@
The Kerr spacetime saturates the inequality (2) with equality for all sub-extremal values $|a|\leq M$:
@@
We compute the right-hand side explicitly. Let $s=\sqrt{M^{2}-a^{2}}$, so that $r_{+}=M+s$.
Step 1: Compute $A/(16\pi)$.
$\frac{A}{16\pi}=\frac{8\pi M(M+s)}{16\pi}=\frac{M(M+s)}{2}.$
Step 2: Compute $4\pi J^{2}/A$.
$\frac{4\pi J^{2}}{A}=\frac{4\pi M^{2}a^{2}}{8\pi M(M+s)}=\frac{Ma^{2}}{2(M+s)}.$
Step 3: Add the terms.
$\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}$ $=\frac{M(M+s)}{2}+\frac{Ma^{2}}{2(M+s)}$ (8)
$=\frac{M(M+s)^{2}+Ma^{2}}{2(M+s)}$ (9)
$=\frac{M[(M+s)^{2}+a^{2}]}{2(M+s)}.$ (10)
Step 4: Simplify $(M+s)^{2}+a^{2}$.
$(M+s)^{2}+a^{2}$ $=M^{2}+2Ms+s^{2}+a^{2}$ (11)
$=M^{2}+2Ms+(M^{2}-a^{2})+a^{2}\quad\text{(since }s^{2}=M^{2}-a^{2}\text{)}$ (12)
$=2M^{2}+2Ms=2M(M+s).$ (13)
Step 5: Final computation.
$\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}=\frac{M\cdot 2M(M+s)}{2(M+s)}=M^{2}.$
Therefore:
$\sqrt{\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}}=M=M_{\mathrm{ADM}}.$
This confirms Kerr saturation with equality. ∎
Corollary 2.4 (Special Cases of Kerr Saturation).
  1. (1)
    Schwarzschild ($a=0$): $A=16\pi M^{2}$, $J=0$, and the bound reduces to $M\geq\sqrt{A/(16\pi)}=M$. ✓
  2. (2)
    Extremal Kerr ($a=M$): $A=8\pi M^{2}$, $J=M^{2}$, giving $\sqrt{A/(16\pi)+4\pi J^{2}/A}=\sqrt{M^{2}/2+M^{2}/2}=M$. ✓
@@
The Kerr initial data $(g_{\mathrm{Kerr}},K_{\mathrm{Kerr}})$ on a Boyer–Lindquist constant-$t$ slice is asymptotically flat and satisfies the standard Kerr falloff. In particular, in the coordinates $(r,\theta,\phi)$ with $r>r_{+}$ (exterior region):
@@
    $g_{ij}-\delta_{ij}=O(M/r)\in C^{k,\beta}_{-1}$ for all $k$;
@@
    $K_{ij}=O(Ma/r^{2})\in C^{k,\beta}_{-2}$ for all $k$;
@@
    The decay rate $\tau=1>1/2$ ensures well-defined ADM mass $M_{\mathrm{ADM}}=M$.
Moreover, by working in horizon-penetrating coordinates (e.g., Kerr–Schild or ingoing Eddington–Finkelstein), one can extend the induced geometry across the horizon (and in particular across the bifurcation sphere) in a smooth fashion; see, for example, standard GR references such as [17].
Thus Kerr data satisfies the hypotheses of Theorem 1.2 in the strictly sub-extremal case $0<|a|<M$. In the extremal limit $|a|=M$, one should treat near-horizon regularity and the precise formulation of “slice through the horizon” with additional care; for this reason we only claim (H1)–(H3) uniformly for the full range $|a|\leq M$.
Example 2.6 (Worked Numerical Example: Near-Extremal Kerr with $a/M=0.9$).
Consider a near-extremal Kerr black hole with spin parameter $a=0.9M$, demonstrating that the bound is saturated for all sub-extremal values.
Step 1: Compute derived quantities.
$s$ $=\sqrt{M^{2}-a^{2}}=\sqrt{M^{2}-0.81M^{2}}=\sqrt{0.19}M\approx 0.4359M,$
$r_{+}$ $=M+s\approx 1.4359M,$
$J$ $=aM=0.9M^{2}.$
Step 2: Compute horizon area.
$A$ $=4\pi(r_{+}^{2}+a^{2})=4\pi[(1.4359M)^{2}+(0.9M)^{2}]$
$=4\pi[2.0618M^{2}+0.81M^{2}]=4\pi\cdot 2.8718M^{2}\approx 11.4872\pi M^{2}.$
Step 3: Verify the bound.
$\frac{A}{16\pi}$ $=\frac{11.4872\pi M^{2}}{16\pi}\approx 0.7180M^{2},$
$\frac{4\pi J^{2}}{A}$ $=\frac{4\pi(0.81M^{4})}{11.4872\pi M^{2}}\approx 0.2820M^{2},$
$\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}$ $\approx 0.7180M^{2}+0.2820M^{2}=1.0000M^{2}.$
Therefore $\sqrt{A/(16\pi)+4\pi J^{2}/A}=M$, confirming exact saturation.
Physical interpretation: As $a/M$ increases from 0 (Schwarzschild) to 1 (extremal), the two terms in the bound exchange dominance. The area term $A/(16\pi M^{2})$ decreases while the angular momentum term $4\pi J^{2}/(AM^{2})$ increases, but their sum remains exactly $M^{2}$:
$a/M$ $A/(16\pi M^{2})$ $4\pi J^{2}/(AM^{2})$
0 (Schwarzschild) 1.000 0.000
0.5 0.933 0.067
0.9 (this example) 0.718 0.282
0.99 0.571 0.429
1.0 (extremal) 0.500 0.500
The sum is always exactly $M^{2}$ for Kerr, confirming saturation across the entire sub-extremal range $|a|\leq M$.
Remark 2.7 (Summary: Kerr Saturation Verified).
Substituting the Kerr horizon parameters ($A=8\pi M(M+\sqrt{M^{2}-a^{2}})$ and $J=aM$) into the inequality yields exactly $M_{\mathrm{ADM}}=M$, confirming the bound is sharp. This verification is a necessary consistency check: if Kerr failed to saturate the bound, the conjecture would be false. The explicit algebraic identity $A/(16\pi)+4\pi J^{2}/A=M^{2}$ for Kerr (Theorem 2.3) determines the correct functional form of the angular momentum Penrose inequality.
3 Proof Strategy
The proof employs a four-stage method extending the techniques of Bray–Khuri [11], Han–Khuri [29], and Agostiniani–Mazzieri–Oronzio [1]. The strategy transforms the initial data $(M,g,K)$ into a manifold with non-negative scalar curvature, on which a monotonicity argument can be applied.
3.1 Overview of the Construction
The proof proceeds through the following stages:
  1. 1.
    Jang Equation: We solve the axisymmetric Jang equation to construct a graph $\Sigma$ in the product manifold $M\times\mathbb{R}$. The resulting Jang manifold $(\bar{M},\bar{g})$ has a cylindrical end corresponding to the original MOTS and satisfies a scalar curvature inequality involving the energy density.
  2. 2.
    Conformal Deformation: We solve a generalized Lichnerowicz equation for a conformal factor $\phi$, producing a metric $\tilde{g}=\phi^{4}\bar{g}$ with non-negative scalar curvature $R_{\tilde{g}}\geq 0$. The source term in this equation involves the angular momentum, ensuring that the geometry encodes the rotation.
  3. 3.
[…]
  4. 4.
    Monotonicity and Boundary Values: We define the AM-Hawking mass $m_{H,J}(t)$ and prove its monotonicity along the flow. The sub-extremality condition $A\geq 8\pi|J|$ ensures the non-negativity of the derivative. Evaluating the mass at the horizon ($t=0$) and at infinity ($t=1$) yields the inequality.
3.2 Comparison with Previous Methods
Our approach differs from the standard Riemannian Penrose inequality proofs [32, 10] by incorporating angular momentum into the flow and the curvature condition. Unlike the inverse mean curvature flow or conformal flow used in the time-symmetric case, we use the $p$-harmonic flow, which allows for a natural coupling with the angular momentum conservation via the level set geometry.
3.3 Technical Components
The proof relies on four main technical results, which we establish in the subsequent sections:
  1. (T1)
    Jang Existence4): The existence of solutions to the axisymmetric Jang equation with controlled blow-up at the MOTS.
  2. (T2)
    AM-Lichnerowicz5): The solvability of the conformal problem with the angular momentum source term.
  3. (T3)
    Conservation6): The conservation of $J$ along the flow.
  4. (T4)
    Sub-Extremality7): The preservation of the sub-extremality condition $A(t)\geq 8\pi|J|$ along the flow.
3.4 Bounded Geometry Verification
A key technical assumption used throughout the proof is “bounded geometry” of the initial data and derived manifolds. We now verify that this assumption is satisfied for initial data in the class considered by Theorem 1.2.
Lemma 3.1 (Bounded Geometry for Axisymmetric Vacuum Data).
Let $(M,g,K)$ be asymptotically flat, axisymmetric, vacuum initial data with decay rate $\tau>1/2$ and outermost strictly stable MOTS $\Sigma$. Then:
  1. (i)
    Curvature bounds: There exist constants $C_{R},C_{K}>0$ depending only on $(M,g,K)$ such that:
    $|\mathrm{Rm}_{g}|\leq C_{R},\quad|\nabla\mathrm{Rm}_{g}|\leq C_{R},\quad|K|\leq C_{K},\quad|\nabla K|\leq C_{K}$
    on any compact subset of $M$.
  2. (ii)
    Injectivity radius: There exists $\iota_{0}>0$ such that $\mathrm{inj}(M,g)\geq\iota_{0}$ on any compact subset bounded away from $\Sigma$.
  3. (iii)
    MOTS geometry bounds: The stable MOTS $\Sigma$ satisfies:
    $|A_{\Sigma}|^{2}\leq C_{A},\quad|\nabla^{\Sigma}A_{\Sigma}|\leq C_{A},\quad\lambda_{1}(L_{\Sigma})\geq\lambda_{0}>0,$
    where $A_{\Sigma}$ is the second fundamental form and $L_{\Sigma}$ is the MOTS stability operator.
@@ 3.4 Bounded Geometry Verification
    Jang manifold bounds: The Jang manifold $(\bar{M},\bar{g})$ from Theorem 4.12 satisfies:
    $|\mathrm{Rm}_{\bar{g}}|\leq C_{\bar{g}},\quad\mathrm{inj}(\bar{M},\bar{g})\geq\iota_{\bar{g}}>0$
    away from the cylindrical end, and the cylindrical end metric satisfies exponential convergence to the product $dt^{2}+g_{\Sigma}$ with rate $\beta_{0}=2\sqrt{\lambda_{1}(L_{\Sigma})}>0$.
@@
    Conformal metric bounds: The conformal metric $\tilde{g}=\phi^{4}\bar{g}$ from Theorem 5.8 satisfies:
    $C^{-1}\bar{g}\leq\tilde{g}\leq C\bar{g},\quad|\mathrm{Rm}_{\tilde{g}}|\leq C_{\tilde{g}}$
    for some $C>1$ depending on the initial data.
@@
(i) Curvature bounds. For asymptotically flat data with decay rate $\tau>1/2$, the constraint equations
$R_{g}=|K|^{2}-(\mathrm{tr}K)^{2}+2\mu,\quad D^{j}K_{ij}-D_{i}(\mathrm{tr}K)=\boldsymbol{j}_{i}$
with $\mu=\boldsymbol{j}=0$ (vacuum) imply that the scalar curvature is determined algebraically by $K$. Since $K_{ij}=O(r^{-\tau-1})$ with bounded derivatives, the Ricci tensor satisfies $\mathrm{Ric}_{g}=O(r^{-2\tau-2})$. By elliptic regularity for the vacuum constraint equations (Bianchi identity), all curvature derivatives are controlled. On any compact set, these bounds are finite.
(ii) Injectivity radius. By the Cheeger–Gromov compactness theorem, manifolds with bounded curvature and positive lower volume bound have positive injectivity radius. For asymptotically flat manifolds, this holds on compact subsets. Near $\Sigma$, the injectivity radius may degenerate, but we work away from $\Sigma$ (or on the Jang manifold where $\Sigma$ is “blown up” to infinity).
(iii) MOTS geometry. For a strictly stable MOTS ($\lambda_{1}(L_{\Sigma})>0$) in vacuum data satisfying DEC:
  • The Galloway–Schoen theorem [27] implies $\Sigma\cong S^{2}$ with positive Gaussian curvature somewhere;
  • Stability bounds the second fundamental form: by the stability inequality $\int_{\Sigma}(|A_{\Sigma}|^{2}+\mathrm{Ric}_{g}(\nu,\nu))\psi^{2}\leq\int_{\Sigma}|\nabla\psi|^{2}$ for the principal eigenfunction $\psi>0$, we have $\|A_{\Sigma}\|_{L^{2}}^{2}\leq C(\lambda_{1},\text{geom})$;
  • Higher regularity follows from elliptic estimates on the MOTS equation $\theta^{+}=0$.
(iv) Jang manifold. The Jang metric $\bar{g}=g+df\otimes df$ differs from $g$ by a rank-1 perturbation. Away from $\Sigma$, where $|\nabla f|$ is bounded, the curvature of $\bar{g}$ is controlled by that of $g$ plus terms involving $\nabla^{2}f$, which are bounded by the Jang equation. Near the cylindrical end, the exponential convergence to the product metric gives explicit bounds. The injectivity radius is positive on any compact subset of $\bar{M}$.
(v) Conformal bounds. The conformal factor $\phi$ from Theorem 5.8 satisfies $0<c_{\phi}\leq\phi\leq C_{\phi}$ (bounded away from 0 and $\infty$) by the maximum principle and asymptotic analysis. The conformal transformation formula
$\mathrm{Rm}_{\tilde{g}}=\phi^{-4}(\mathrm{Rm}_{\bar{g}}-2\phi^{-1}\nabla^{2}_{\bar{g}}\phi+\text{lower order})$
then gives curvature bounds for $\tilde{g}$ in terms of those for $\bar{g}$ and the $C^{2}$ norm of $\phi$. ∎
Remark 3.2 (Uniformity of Constants).
The constants in Lemma 3.1 depend on the initial data $(M,g,K)$ but are finite and computable for any data in the class of Theorem 1.2. In particular:
  • The “bounded geometry” assumptions used in estimates throughout this paper (e.g., in Lemma 6.39, Remark 4.17, and the Willmore derivative bound in (75)) are verified for our class of data by Lemma 3.1.
  • The proof does not require any “generic” assumptions beyond those stated in Theorem 1.2.

### v4

@@ 1.3 Relation to Prior Work
The Penrose inequality with angular momentum has a substantial history. Our proof synthesizes the Jang equation methods of Bray–Khuri [11] and Han–Khuri [29], the area-angular momentum bounds of Dain–Reiris [22], and the $p$-harmonic flows of Agostiniani–Mazzieri–Oronzio [1].
The time-symmetric case ($J=0$) was established by Huisken–Ilmanen [32] using inverse mean curvature flow and by Bray [10] via conformal flow. The Jang equation approach to the non-time-symmetric case, along with the scalar curvature identity on the Jang manifold, was developed by Bray–Khuri; Han–Khuri refined the blow-up analysis at the MOTS. The sub-extremality bound $A\geq 8\pi|J|$ for stable axisymmetric MOTS is due to Dain–Reiris, and we use it as an input rather than reproving it. The $p$-harmonic flow machinery and Hawking mass monotonicity follow Agostiniani–Mazzieri–Oronzio; our contribution is adapting their framework to track angular momentum and establishing the conservation theorem.
Our main new ingredients are: (1) the modified Hawking mass $m_{H,J}(t)=\sqrt{m_{H}^{2}+4\pi J^{2}/A(t)}$ and its monotonicity; (2) angular momentum conservation along the flow under the vacuum hypothesis; (3) the extension of the Jang analysis to incorporate twist.
@@ 1.4 Organization
Section 2 verifies the inequality for Kerr. Section 3 outlines the proof strategy. Sections 47 develop the four technical stages: the Jang equation, the Lichnerowicz equation, angular momentum conservation, and sub-extremality preservation. Section 8 assembles these into the proof of Theorem 1.1. Section 9 treats the equality case, and Section 10 discusses open problems.
@@ 2 Verification for Kerr Spacetime
Before proceeding to the proof, we verify that Kerr spacetime saturates the inequality. This is a necessary consistency check.
Working in geometric units ($G=c=1$), the Kerr spacetime with mass $M$ and spin parameter $a=J/M$ (where $|a|\leq M$) has outer horizon radius $r_{+}=M+\sqrt{M^{2}-a^{2}}$ and horizon area
$A=4\pi(r_{+}^{2}+a^{2})=8\pi M(M+\sqrt{M^{2}-a^{2}}).$
@@
For all sub-extremal Kerr black holes ($|a|\leq M$),
@@
Set $s=\sqrt{M^{2}-a^{2}}$, so $r_{+}=M+s$. Then
$\frac{A}{16\pi}=\frac{M(M+s)}{2},\qquad\frac{4\pi J^{2}}{A}=\frac{Ma^{2}}{2(M+s)}.$
Adding these gives
$\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}=\frac{M[(M+s)^{2}+a^{2}]}{2(M+s)}.$
Since $s^{2}=M^{2}-a^{2}$, we have $(M+s)^{2}+a^{2}=2M(M+s)$, and so
$\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}=M^{2}.\qed$
As special cases: for Schwarzschild ($a=0$), we have $A=16\pi M^{2}$ and $J=0$, giving the standard bound; for extremal Kerr ($a=M$), we have $A=8\pi M^{2}$ and $J=M^{2}$, and the two terms contribute equally to give $M$.
@@
On a Boyer–Lindquist slice with $r>r_{+}$, Kerr initial data satisfies $g_{ij}-\delta_{ij}=O(M/r)$ and $K_{ij}=O(Ma/r^{2})$, with decay rate $\tau=1>1/2$. The data satisfies hypotheses (H1)–(H4) for strictly sub-extremal parameters; see [17].
3 Proof Strategy
The proof extends the Jang equation approach of Bray–Khuri [11] and Han–Khuri [29], combined with the $p$-harmonic monotonicity of Agostiniani–Mazzieri–Oronzio [1]. The idea is to transform the initial data into a manifold with non-negative scalar curvature and then apply a flow argument.
3.1 Outline
The argument has four stages:
(i) Jang equation. We solve the axisymmetric Jang equation to obtain a graph in $M\times\mathbb{R}$. The Jang manifold $(\bar{M},\bar{g})$ has a cylindrical end at the MOTS and satisfies a scalar curvature inequality from the Bray–Khuri identity.
(ii) Conformal deformation. We solve a Lichnerowicz-type equation for a conformal factor $\phi$ to produce $\tilde{g}=\phi^{4}\bar{g}$ with $R_{\tilde{g}}\geq 0$.
[…]
(iv) Monotonicity. We define a modified Hawking mass $m_{H,J}(t)$ which incorporates angular momentum. This mass is monotone increasing along the flow; evaluating at the boundary gives the inequality.
3.2 Technical Results
Details appear in §4 (Jang equation), §5 (Lichnerowicz equation), §6 (angular momentum conservation), and §7 (sub-extremality).
3.3 Bounded Geometry
The following lemma collects the geometric bounds used throughout.
Lemma 3.1.
Let $(M,g,K)$ satisfy the hypotheses of Theorem 1.1. Then:
@@ 3.3 Bounded Geometry
    Curvature and its derivatives are bounded on compact subsets of $M$.
@@
    The injectivity radius is positive on compact subsets away from $\Sigma$.
@@
    The MOTS $\Sigma$ has bounded second fundamental form and $\lambda_{1}(L_{\Sigma})>0$.
@@
    The Jang manifold $(\bar{M},\bar{g})$ has bounded curvature away from the cylindrical end, where it converges exponentially to a product.
@@
    The conformal metric $\tilde{g}=\phi^{4}\bar{g}$ satisfies $C^{-1}\bar{g}\leq\tilde{g}\leq C\bar{g}$ for some $C>1$.
@@
Part (i) follows from the vacuum constraints and asymptotic flatness. Part (ii) is Cheeger–Gromov. For (iii), the Galloway–Schoen theorem [27] gives spherical topology, and stability bounds $|A_{\Sigma}|$ via the stability inequality. Parts (iv) and (v) follow from the Jang equation and maximum principle arguments for $\phi$; see the constructions in §4 and §5. ∎